# What Happens During the Loss Plateau? Understanding Abrupt Learning in Transformers

### v1

@@ 5 A Further Look at Repetition Bias
5.1 Beyond Repetitions in Consecutive Tokens
@@
${\rm SeqEnt}(y_{1},\ldots,y_{n}):=\sum_{i=1}^{\left|V\right|}p_{i}\log(1/p_{i});\quad p_{i}=\frac{\left|\{y_{j}=v_{i},j\in[n]\}\right|}{n}$
i.e. simply the entropy of the empirical distribution of tokens in the sequence. Intuitively, the entropy is lower if most probability mass is concentrated at a few tokens, and larger if the tokens are more uniformly distributed. We find that the model output entropy quickly goes to quite low values early in training compared to the entropy of ground-truth data (Figure 6), indicating that the model still has a form of repetition bias. Further, representation collapse still happens in the early phase, with the average cosine similarity going to $0.8$ during the plateau.
@@
Hence, we find that repetition bias might take different forms depending on the task, but still robustly occurs in the early phase of training.
5.2 Repetitive Sequences Are Easier to Learn
@@
On the other hand, we study what happens when the ground-truth data have a lot of repetitions. We consider a simple task REPEAT1 of the form $x_{1},x_{2},\ldots,x_{n},{\rm SEP},y_{1},y_{2},\ldots,y_{n}$, where $y_{i}=x_{1}\,\,\forall i$. Unlike other tasks, the loss curve for REPEAT1 does not have any noticeable plateau, though the accuracy still shows a small plateau period (Figure 7). This observation indicates that such repetitive sequences are easier from an optimization perspective and hence likely “preferred” during the early stage of training. In fact, just one gradient step is sufficient to bring the average representation cosine similarity to $\approx 0.5.$ We show similar results for other task variants REPEAT2, REPEAT4 in Appendix C.
@@ 6 Repetition Bias and Representation Collapse in LLMs
Having established that the degenerate patterns of repetition bias and representation collapse are prevalent in small Transformers trained on algorithmic tasks, a crucial question is whether these phenomena also happen beyond toy settings, specifically during the early pre-training stages of LLMs. We verify that this is indeed the case, using early-stage checkpoints of open-source LLMs Pythia biderman2023pythia and OLMo-2 olmo20242olmo2furious .
@@
Similar representation collapse patterns as Pythia are observed for the OLMo-2 7B model. For its earliest available training checkpoint (step 150, OLMo-2-1124-7B), the average representation cosine similarity in the setup from Section 6 is $\approx 0.93$; for the next checkpoint at step 600, this value has already decreased to $\approx 0.43$ (similar for both greedy decoding and random sampling strategies).
Hence, repetition bias and representation collapse occur in the early pre-training phase of LLMs, validating our findings beyond toy settings.
Figure 8: Representation Collapse during Pythia Pretraining. Representation collapse at different Pythia pretraining checkpoints; tokens generated via greedy decoding.
@@ 7 Related Work
Abrupt learning has been studied in multiple settings; barak2022hidden studied it for parity tasks for multiple neural net architectures, while edelman_evolution studied abrupt learning for a Markov chain task with Transformers. gopalani2024abrupt showed that training a BERT model on matrix completion leads to abrupt learning, while percolation connected abrupt learning for a grammar data setup to graph percolation. For abrupt learning in in-context learning garg2023transformerslearnincontextcase , there has been a line of recent works zhang2025incontext ; singh2024 ; wang2025rich ; chen2024trainingdynamics ; reddy2023mechanisticbasisdatadependence ; zucchet2025emergencesparseattentionimpact that proposed various theoretical and empirical explanations. We aim to understand a unifying reason behind such observations, in an algorithmic setup for multiple tasks with multi-token output sequences, without restrictive assumptions on our model or training setup.
A line of recent work focused on understanding a related phenomenon of grokking power2022grokking , which is abrupt generalization after an extended phase of memorization of training data by the model. Multiple works have studied grokking from the perspective of circuits nanda2023progress ; varma_circuit ; merrill_circuit , representation learning liu_rep_grokking , delay in feature learning kumar2024grokkingtransitionlazyrich ; lyu2024dichotomyearlylatephase ; xu2023benignoverfittinggrokkingrelu , learning syntax for linguistic data chen2024sudden , and specific properties of softmax attention eureka ; prieto2025 . We note that grokking is a fixed-dataset phenomenon, while abrupt learning studied in our work is for online training (‘infinite-data’), hence there is no notion of training data memorization. Saddle-to-saddle dynamics gradualrank have been used to theoretically explain plateaus and sudden drops in loss during training transformers; however, such results require assumptions such as diagonal attention and very small scale of initialization ($\rightarrow 0$). Ideas from statistical physics have also been used towards understanding initial loss plateaus in neural net training saad_solla ; Inoue_2003 ; they work in a 2-layer teacher-student neural net setup, where the second layer is fixed during training, and use order parameters to study training. They show that there is a permutation symmetry in the weight vectors of the first layer during the early plateau stage, and exiting this symmetry state is what leads to drop in loss. Another line of work towards understanding stagewise training dynamics of Transformers uses techniques from singular learning theory hoogland2025losslandscapedegeneracydrives ; the core idea is to estimate the Local Learning Coefficient (LLC) during training, and use this quantity to explain degeneracy in the loss landscape, and consequently the stage-wise training dynamics of Transformers. The interplay of simplicity bias and Transformer learning dynamics has also been studied recently in rende2025distributionalsimplicitybiaslearning ; belrose2024neuralnetworkslearnstatistics . In belrose2024neuralnetworkslearnstatistics authors show that neural nets learn to use lower-order moments of data earlier in training, and that the embedding statistics of Transformer models and token $n$-gram frequencies are related, explaining a specific distributional simplicity bias during training. rende2025distributionalsimplicitybiaslearning show that Transformer-based models progressively learn higher-order (‘many-body’) interactions between tokens in the sequence.
@@
Repetition in language model outputs is a well studied problem holtzman2020curiouscaseneuraltext ; hiraoka2025repetitionneuronslanguagemodels ; li2023repetition ; fu2021theoreticalanalysisrepetitionproblem ; yao2025understandingrepeatcurselarge ; wang2024mitigatinglanguagemismatchrepetition ; xu2022learningbreakloopanalyzing ; ildiz2024selfattentionmarkovmodelsunveiling . However these works focus not on the early phase of training, but on how repetition may arise in pretrained models, and how to mitigate such phenomena. choshen2022grammarlearningtrajectoriesneurallanguage remarked that in the early phase of training language models, the output might contain some word repetitions, but understanding this occurrence is not the main focus of their work. Rank collapse is a related phenomenon for deep softmax transformers at initialization that might hinder training anagnostidis2022signal ; however, our representation collapse phenomenon is different in that (i) we use shallow (1 or 2 layers) Transformers instead of deep ones; (ii) we use linear attention instead of softmax; (iii) our observed representation collapse occurs only after a few steps of training, not at initialization. Note that we study representation collapse in the early phase of training, which is distinct from the notion of representation collapse in barbero2024transformersneedglassesinformation ; they show that for 2 sequences $(v_{1},v_{2},\ldots,v_{n})$ and $(v_{1},v_{2},\ldots,v_{n},v_{n})$, as $n$ grows large, the pretrained model’s hidden state representation for the last token becomes identical for both sequences (Theorem 4.2, barbero2024transformersneedglassesinformation ).
@@ 8 Discussion
We identified repetition bias and representation collapse as key characteristics of the early-phase implicit biases of Transformer training, which are closely connected to the commonly observed loss plateau. The question of why such degeneracy exists during early time training is an important question for future work. Furthermore, while we showed that the slow rate of learning attention map leads to model remaining in the plateau for an extended time, why the rate of learning attention map is slow for algorithmic tasks, and how it connects to the intuitive “complexity” of the task is an interesting question.
Acknowledgements
PG would like to thank Misha Belkin, Ekdeep Singh Lubana, Yongyi Yang, Zhiwei Xu and Vaibhav Balloli for useful discussions at various stages of this work. Part of this work was done when the authors were visiting the Modern Paradigms in Generalization Program at the Simons Institute for the Theory of Computing.
@@ Appendix A Implementation Details
Linear attention transformer is obtained simply by removing the softmax activation function when computing the attention map, and setting the causal mask to 00 instead of $\infty$- ∞. We use the existing minGPT implementation [22] (MIT licence) for our experiments, modifying the code as above and wherever required.
@@
We use Pythia [5] / OLMo-2 [31] pretrained models (Apache 2.0 Licence) hosted on Huggingface Transformers [42] and run them on the ARC-Easy dataset [10] (CC-BY-SA 4.0 Licence). We set the use_cache=False in the generate function, and use the hidden state used for predicting each of the 8 output tokens. For random sampling, we use do_sample=True (using default temperature value), using do_sample=False for our greedy decoding results.
@@ Appendix B Results for Other Algorithmic Tasks
Prefix Sum (PRE) Compute prefix sum of a given $n$–length sequence First input element
@@
Multi-Digit Addition (ADD) Add atmost–$n$–digit numbers First digit (00 or $1$) i.e. total carry-over from $n$ digits
Histogram (HIST) Compute counts of each element in $n$–length sequence $\approx 100\%$ Repetitive sequences
Reverse (REV) Reverse $n$–length input sequence Repetitive sequences1
Copy (COPY) Copy $n$–length input sequence Repetitive sequences1
@@ B.1 Multi-Digit Addition
This task involves adding 2 atmost $4$–digit numbers; if the numbers are represented as $a=\overline{a_{1}a_{2}a_{3}a_{4}},\,\,b=\overline{b_{1}b_{2}b_{3}b_{4}}$ and their sum $a+b=c=\overline{c_{0}c_{1}c_{2}c_{3}c_{4}}$ then the training sequences for ADD are of the form
@@
Note that the output sequence is reversed, following the observations from [24]. We find similar abrupt learning characteristics (Figure 9), partial solution in this case being $c_{0}$ i.e. total carry-over from $4$ single digit add operations. An interpretable attention map learnt for the output sequence is shown in Figure 10.
@@ B.2 Prefix sum
[…]
Training dynamics for this task are shown in Fig. 11 which show similar abrupt learning behavior as MWS and partial solution learning for $y_{1}$. The interpretable attention map learnt for this task is shown in Figure 12.
@@ B.3 Permutation
[…]
@@
where $x_{i}\sim{\rm Unif}\{17,18,\ldots,32\},n=16,{\rm SEP}=0$. The partial solution in this case is the output sequence being an permutation of the input sequence $x_{1},\ldots,x_{n}$ i.e., it learns to copy the tokens correctly, but in wrong order (Figure 13). The interpretable attention maps in this case show that the model learns to copy the correct tokens based on the permutation provided (Figure 14(a)) and then uses them for the final output sequence (Figure 14(b)).
@@ B.4 Histogram
where $x_{i}\sim{\rm Unif}\{1,2,\ldots,12\},n=16,{\rm SEP}=0$. We train a 2-layer, 1-head transformer for this task, with gradient clipping $(1.0)$ to avoid loss spikes (Figure 15). We note that the repetition bias in this case is quite strong which leads to $\approx 100\%$ repetitions in the early phase of training, and which we characterize as partial solution for this task. Further we only consider the attention map from layer 1 Figure 16 since this is the most consistent and clearly interpretable across runs, and indicates an identity-map-like function.
Figure 15: Training dynamics for Histogram task. (left) Train/Test Loss and Accuracy; (middle) Repetition frequency and representation collapse; (right) Attention progress measure. We only measure attention progress for the 1st layer, since that is the one that consistently and clearly shows an interpretable pattern (Figure 16).
@@ B.5 Reverse
for $x_{i}\sim{\rm Unif}\{1,2,\ldots,16\},n=16,{\rm SEP}=0$. The training dynamics are shown in Figure 17(a) and the interpretable attention map is shown in Figure 17(b).
@@ B.6 Copy
for $x_{i}\sim{\rm Unif}\{1,2,\ldots,16\},n=16,{\rm SEP}=0$. The training dynamics are shown in Figure 18(a) and the interpretable attention map is shown in Figure 18(b).
@@
Appendix C Results for REPEAT2 and REPEAT4

### v2

@@ 5 A Further Look at Repetition Bias
Beyond Repetitions in Consecutive Tokens
@@
${\rm SeqEnt}(y_{1},\ldots,y_{n}):=\sum_{i=1}^{\left|V\right|}p_{i}\log(1/p_{i});\quad p_{i}=\frac{\left|\{y_{j}=v_{i},j\in[n]\}\right|}{n}$ (6)
i.e. simply the entropy of the empirical distribution of tokens in the sequence. Intuitively, the entropy is lower if most probability mass is concentrated at a few tokens, and larger if the tokens are more uniformly distributed. We find that the model output entropy quickly goes to quite low values early in training compared to the entropy of ground-truth data (Figure˜6), indicating that the model still has a form of repetition bias. Further, representation collapse still happens in the early phase, with the average cosine similarity going to $0.8$ during the plateau. Hence, we find that repetition bias might take different forms depending on the task, but still robustly occurs in the early phase of training.
@@
Repetitive Sequences are Easier to Learn
@@
We study what happens when training our model on data that has a lot of repetitions. We consider a simple task REPEAT1 of the form $x_{1},x_{2},\ldots,x_{n},{\rm SEP},y_{1},y_{2},\ldots,y_{n}$, where $y_{i}=x_{1}\,\,\forall i$. Unlike other tasks, the loss curve for REPEAT1 does not have any noticeable plateau, though the accuracy still shows a small plateau period (Figure˜7). This observation indicates that such repetitive sequences are easier from an optimization perspective and hence likely “preferred” during the early stage of training. In fact, just one gradient step is sufficient to bring the average representation cosine similarity to $\approx 0.5.$ We show similar results for other task variants (REPEAT2, REPEAT4) in Appendix˜D.
@@ 6 Repetition Bias and Representation Collapse in LLMs
Having established that the degenerate patterns of repetition bias and representation collapse are prevalent in small Transformers trained on algorithmic tasks, a crucial question is whether these phenomena also happen beyond toy settings, specifically during the early pre-training stages of LLMs. We verify that this is indeed the case, using early-stage checkpoints of open-source LLMs Pythia [6] and OLMo-2 [37].
@@
Similar representation collapse patterns as Pythia are observed for the OLMo-2 7B model. For its earliest available training checkpoint (step 150, OLMo-2-1124-7B), the average representation cosine similarity in the setup from Section˜6 is $\approx 0.93$; for the next checkpoint at step 600, this value has already decreased to $\approx 0.43$ (similar for both greedy decoding and random sampling strategies). Hence, repetition bias and representation collapse occur in the early pre-training phase of LLMs, validating our findings beyond toy settings.
Figure 8: Representation Collapse, Repetitions during Pythia Pretraining. Representation collapse at different Pythia pretraining checkpoints evaluated on the ARC-Easy test dataset; tokens generated via greedy decoding. We find that similar to our toy experiments, average pairwise cosine similarity between hidden states and repetition frequency start at relatively low values, and increase to near $1.0$ during the early phase of pretraining, followed by a decrease around step $1000.$
@@ 7 Related Work
Abrupt learning has been studied in various settings; [9] study sudden drop in loss when training BERT, and show abrupt learning of the attention map (termed Syntactic Attention Structure / SAS) concurrent with the sudden drop in training loss. [20] study sudden drop in loss (‘Eureka moments’) when training Transformers on multi-step tasks. They show that ill-distributed softmax attention scores lead to small gradients for key, query weights causing slow learning of attention map, and that appropriately modifying softmax temperature can alleviate optimization issues. [3] studied abrupt learning in various neural net architectures trained on sparse parity tasks, demonstrating underlying hidden progress during training via measures such as Fourier gap and weight movement. [18] study abrupt learning when training a BERT model on matrix completion, while [32] analyze abrupt learning for a grammar data setup through graph percolation. For abrupt learning in in-context learning [17], there has been a line of recent works [54, 44, 47, 10, 41, 55, 15] that proposed various theoretical and empirical explanations. [15] studied abrupt learning for a Markov chain in-context learning task with Transformers, and demonstrate a partial solution in their setup in the form of a unigram estimator. [46] analyze partial solutions for an in-context nn–gram task with Transformers, corresponding to kk–gram estimators $(k<n)$. For diagonal-attention Transformers with small initialization, [7] analyze ‘saddle-to-saddle’ dynamics for a theoretical understanding of loss plateaus during training.
A line of recent work has focused on understanding ‘grokking’ [39], i.e. abrupt generalization after an extended phase of memorization of training data by the model. Subsequent works have studied grokking through circuit formation [36, 45, 34], representation learning [31], delay in feature learning [27, 33, 52]. [40] further study ‘naive loss minimization’ and numerical stability issues for understanding grokking. Note that grokking is a fixed-dataset phenomenon concerning memorization of the training dataset, and hence somewhat distinct from abrupt learning studied in this paper.
@@
Rank collapse is a related phenomenon for deep softmax transformers at initialization that might hinder training [2]; however, our representation collapse phenomenon is different in that (i) we use shallow (1 or 2 layers) Transformers instead of deep ones; (ii) we use linear attention instead of softmax; (iii) our observed representation collapse occurs only after a few steps of training, not at initialization. [4] show a form of representation collapse so that for 2 sequences $(v_{1},v_{2},\ldots,v_{n})$ and $(v_{1},v_{2},\ldots,v_{n},v_{n})$, as $n$ grows large, the pretrained model’s hidden state representation for the last token becomes identical for both sequences (Theorem 4.2, [4]). Note that we study evolution of representation collapse in the early phase of Transformer training, distinct from the notion of representation collapse at initialization, or in final pretrained models.
Repetition in language model outputs is a well studied problem [21, 19, 30, 16, 53, 48, 51, 23]. However most works focus not on the early phase of training, but on how repetition may arise in trained language models, and how to mitigate them. Towards understanding learning dynamics, [11, 8] report token repetitions in the early phase of training language models. Please see Appendix˜I for additional discussion on related work.
@@ 8 Discussion
We identified repetition bias and representation collapse as key characteristics of the early-phase implicit biases of Transformer training, which are closely connected to the commonly observed loss plateau. We further discussed ‘search’ over sequence tokens as a possible cause for loss plateau. The question of why behaviors such as representation collapse exist during early time training is an important question for future work. Furthermore, while we hypothesize a search-like phenomenon to be the reason behind slow learning of attention map, a rigorous theoretical validation of this hypothesis, and how it possibly connects to the intuitive “complexity” of the task, are important questions for further research.
Acknowledgments
This work was supported by DARPA. The authors thank Misha Belkin, Ekdeep Singh Lubana, Yongyi Yang, Zhiwei Xu, Vaibhav Balloli, and anonymous reviewers for helpful comments at various stages of this work. Part of this work was done when the authors were visiting the Modern Paradigms in Generalization Program at the Simons Institute for the Theory of Computing.
@@ Appendix A Implementation Details
Linear attention transformer is obtained simply by removing the softmax activation function when computing the attention map, and setting the causal mask to 0 instead of -$\infty$. We use the existing minGPT implementation [26] (MIT licence) for our experiments, modifying the code as above and wherever required.
@@
We use Pythia [6] / OLMo-2 [37] pretrained models (Apache 2.0 Licence) hosted on Huggingface Transformers [50] and evaluate them on the ARC-Easy / Challenge datasets [12] (CC-BY-SA 4.0 Licence), and GSM8K [13] (MIT Licence). We set the use_cache=False in the generate function, and use the hidden state used for predicting each of the 8 output tokens. For random sampling, we use do_sample=True (using default temperature value), using do_sample=False for our greedy decoding results.
@@ Appendix B Results for Other Algorithmic Tasks
Prefix Sum (PRE) Compute prefix sum of a given nn–length sequence First input element
@@
Multi-Digit Addition (ADD) Add atmost–nn–digit numbers First digit (0 or $1$) i.e. total carry-over from $n$ digits
Histogram (HIST) Compute counts of each element in nn–length sequence $\approx 100\%$ Repetitive sequences
Reverse (REV) Reverse nn–length input sequence Repetitive sequences1
Copy (COPY) Copy nn–length input sequence Repetitive sequences1
Copy + MWS + MWS3 Concatenation of sequences from COPY, MWS, and MWS3 (moving window of 3 elements) During 1st plateau: COPY part correct; 2nd plateau: COPY and MWS parts correct (Section˜B.7)
@@
In the following table we describe APM sets $\Omega$ for different algorithmic tasks used in this work. We assume the attention map is of the shape $A\in R^{(L-1)\times(L-1)}$ in the next token prediction setup on the full input sequence. Hence, for MWS, Prefix sum, Histogram, Reverse, Copy, $L=33$, for permutation $L=50$, and for Multi-digit addition $L=15$. The APM sets are denoted by sets of tuples $(i,j)\in[L-1]\times[L-1]$ indicating the row/column indices in the attention map $A$.
Table 2: Attention Progress Measure (APM) sets $(\Omega)$ for all tasks
Task APM set $\Omega$
MWS $\{(16,0)\}\cup\{(16+i,i-1)\mid i\in[1,15]\}\cup\{(16+i,i)\mid i\in[1,15]\}$
Prefix Sum $\{(16,0)\}\cup\{(16+i,16+i)\mid i\in[1,15]\}\cup\{(16+i,i)\mid i\in[1,15]\}$
Multi-digit Addition $\{(i,12-i)\mid i\in[9,12]\}\cup\{(i,17-i)\mid i\in[9,12]\}\cup\newline \{(i,13-i)\mid i\in[9,13]\}\cup\{(i,18-i)\mid i\in[9,13]\}$
Permutation Layer 1: $\{(17+i,\pi_{i+1}-1)\mid i\in[0,15]\}$
Layer 2: $\{(33+i,17+i)\mid i\in[0,15]\}$
Histogram Layer 1: $\{(16+i,i)\mid i\in[0,15]\}$
Copy $\{(16+i,i)\mid i\in[0,15]\}$
Reverse $\{(16+i,15-i)\mid i\in[0,15]\}$
@@ B.1 Multi-Digit Addition
This task involves adding 2 atmost 44–digit numbers; if the numbers are represented as $a=\overline{a_{1}a_{2}a_{3}a_{4}},\,\,b=\overline{b_{1}b_{2}b_{3}b_{4}}$ and their sum $a+b=c=\overline{c_{0}c_{1}c_{2}c_{3}c_{4}}$ then the training sequences for ADD are of the form
@@
Note that the output sequence is reversed, following the observations from [28]. We find similar abrupt learning characteristics (Figure˜9), partial solution in this case being $c_{0}$ i.e. total carry-over from $4$ single digit add operations. An interpretable attention map learnt for the output sequence is shown in Figure˜10.
@@ B.2 Prefix sum
$y_{i}=\left(\sum_{j=1}^{i}x_{j}\right)\,{\rm mod}\,\,17\quad\forall i\in[n]$
Training dynamics for this task are shown in Fig. 11 which show similar abrupt learning behavior as MWS and partial solution learning for $y_{1}$. The interpretable attention map learnt for this task is shown in Figure˜12.
@@ B.3 Permutation
This task involves training a 2-layer, 1-head Transformer on permuting a lengthn-n sequence using the permutation $\pi$, which is generated at random and is distinct for each training sequence. Formally, for a sequence of positive integers $(x_{1},\ldots,x_{n})$ and a permutation $(\pi_{1},\ldots,\pi_{n})$ over $[n],$ training sequences for PER, $k=0,1,2,\ldots$ are given by
@@
where $x_{i}\sim{\rm Unif}\{17,18,\ldots,32\},n=16,{\rm SEP}=0$. The partial solution in this case is the output sequence being an permutation of the input sequence $x_{1},\ldots,x_{n}$ i.e., it learns to copy the tokens correctly, but in wrong order (Figure˜13). The interpretable attention maps in this case show that the model learns to copy the correct tokens based on the permutation provided (Figure˜14(a)) and then uses them for the final output sequence (Figure˜14(b)).
@@ B.4 Histogram
where $x_{i}\sim{\rm Unif}\{1,2,\ldots,12\},n=16,{\rm SEP}=0$. We train a 2-layer, 1-head transformer for this task, with gradient clipping $(1.0)$ to avoid loss spikes (Figure˜15). We note that the repetition bias in this case is quite strong which leads to $\approx 100\%$ repetitions in the early phase of training, and which we characterize as partial solution for this task. Further we only consider the attention map from layer 1 (Figure˜16) since this is the most consistent and clearly interpretable across runs, and indicates an identity-map-like function.
Figure 15: Training dynamics for Histogram task. (left) Train/Test Loss and Accuracy; (middle) Repetition frequency and representation collapse; (right) Attention progress measure. We only measure attention progress for the 1st layer, since that is the one that consistently and clearly shows an interpretable pattern (Figure˜16).
@@ B.5 Reverse
for $x_{i}\sim{\rm Unif}\{1,2,\ldots,16\},n=16,{\rm SEP}=0$. The training dynamics are shown in Figure˜17(a) and the interpretable attention map is shown in Figure˜17(b).
@@ B.6 Copy
for $x_{i}\sim{\rm Unif}\{1,2,\ldots,16\},n=16,{\rm SEP}=0$. The training dynamics are shown in Figure˜18(a) and the interpretable attention map is shown in Figure˜18(b).
@@
B.7 Concatenation of Copy, MWS, MWS3
We train the model on following sequences,
$x_{1},x_{2},\ldots,x_{16},{\rm SEP},y_{1},y_{2},\ldots,y_{13}$
[…]
with $x_{i}\sim{\rm Unif}\{1,2,\ldots,16\},{\rm SEP}=17,$ and find that the train loss follows a ‘staircase’-like evolution, learning the 3 distinct parts $([y_{1},\ldots,y_{4}],[y_{5},\ldots,y_{9}],[y_{10},\ldots,y_{13}])$ sequentially (Figure˜19). Similar sequential trend is seen for accuracy and average cosine similarity, measured separately over these parts (labeled Part 1,2,3).
Figure 19: Partial Solution for Concatenated Target Sequence.
Appendix C Probing Attention Head Output
We probe the output hidden states of the Attention block, before they are added to the residual stream (Equation˜2). Specifically, we use a 2-layer MLP probe with hidden layer width $256$ and ReLU activation function to map these hidden states to the ground truth output labels $y_{i}$. Formally, for hidden states $\mathbf{h}\in\mathbb{R}^{256}$, the probe is defined as ${\rm Probe}(\mathbf{h}):=W_{P,2}{\rm ReLU}(W_{P,1}\mathbf{h}),\,W_{P,2}\in\mathbb{R}^{17\times 256},W_{P,1}\in\mathbb{R}^{256\times 256},$ implemented using scikit-learn MLPClassifier [35]. We probe at every $10$ training steps of the Transformer model, training the probe on $1024$ samples and evaluating using $1024$ held-out samples, both different from training data for the Transformer.
We find that the probe test accuracy increases earlier than the sudden jump in model accuracy (Figure˜20), demonstrating ‘hidden’ progress in Attention block outputs.
Figure 20: Probing Attention Outputs. We find that the probe accuracy when probing attention block outputs (Equation˜2) starts to increase earlier than the model accruacy during training, highlighting ‘hidden’ gradual progress in the hidden states in addition to attention map (via attention progress measure).
Appendix D Results for REPEAT2, REPEAT4